FEL Upgrade Project Weekly Brief August 26-30, 2002

 

 

Highlights:

We are commencing the Accelerator Readiness Review Process in anticipation of turning on the laser. See the web link below.

There has been much progress in installation around the machine and fabrication of magnets and vacuum components (although much remains). The 3F region of beamline quadrupole girders (downstream of the third cryomodule before the bend) was connected by the vacuum group.

Management:

George Neil attended a Pulsed Laser Lethality Workshop and presented a description of our Upgrade and schedule to allow interested parties to plan on utilizing the laser for materials tests. Discussions were also held with a group from LANL concerning the possibility of testing predictions that at sufficient peak intensities compensation of thermal blooming can be accomplished by using non-linear effects of the atmosphere. It appears the IR Upgrade is in the proper parameter range to test these ideas.

Following the Workshop a separate meeting of the JTO HEL Lethality TAWG was held. Discussions on prioritization of tasks for the coming years included establishing a database of prior measurements and (on the Navy side) planning for establishment of test facility at JLab. A report detailing this roadmap is in preparation for November to present to the JTO.

The internal JLab web-based FEL ARR (Accelerator Readiness Review) process is up an running. Sub-systems managers are using it to document their sub-system's Readiness status and to build an ARR reference documents database. The following is a link to the main ARR Systems Status Home Page. (Note: For security purposes, the link and database is behind the Jlab firewall.) http://devsrv-e.acc.jlab.org/safety/fel/index.php

WBS 3 (Beam Physics):

In response to an inquiry by S. Benson, and with information he kindly provided, a driver configuration supporting the 1 micron FEL option was developed. It is identical to the rev113 design save for an inward move (toward the wiggler) of the quad triplets adjacent to the FEL insertion. This was required to avoid generating large beam envelopes (with associated aberrations) where the small, strongly divergent beam enters/exits the shorter wiggler.

 

 

WBS 4 (Injector):

The support tube was e-beam welded to its flange. The resistor drawings are in sign off. The rear stalk bellows drive unit design is in review. The gun oven extension drawings are in review. All off the internal hardware for assembly of the gun has been vented and is out to be silver plated. Drive hardware for the running resistor is on order.

Gun HVPS - The SF6 piping is nearly complete. The 3 stacks were positioned and the final positioning and pining was done this week. The parts for the interlocks were also located. A system print has been started.

WBS 5 (SRF):

Preparations continue and are nearly complete for measurements of a full cavity HOM damping Qs including high power coupler effects. Modifications to the vertical dewar mounting were readied. In addition estimates of beam breakup thresholds were calculated based on earlier measured values as well as copper model measurements.

WBS 6 (RF):

Quarter HVPS - Work continued this week on the Crowbar and grounding system for the HVPS. The thyratron for the Unit 4 was tested and found to be faulty. The 2 old faulty thyratrons are being returned to the vendor for estimates to restore them to working order. A P.O. has

been placed to replace the thyratron we borrowed and tested from the SNS project. Progress is being made on the grounding system, but 1.5k amps with nanosecond rise time does make healthy transients.

Quarter Klystrons - Hoses were fabricated this week for the Circulator, RF Load, and HOM LCW manifolds.

Zone 4 - The missing waveguides are being fabricated in the Machine Shop and are due next week. The flexible waveguide couplers are also due in next week. They will be assembled when the parts are received.

Zone 1 - J. Preble is scheduling the waveguide connections to the Quarter.

WBS 8 (Instrumentation):

Beam Loss Monitors - the F0151 PCB assembly continues to make progress with EES. The first board is expected to be finished early next week and will immediately be put through testing. The F0150 PCB ECO has been finished and the pcbs are being made new.

Drive Laser Pules Control - The pulse rep-rate and pulse width sections of the control screen is working well. No more progress will be made on the EPICS screen until the next stage of hardware is implemented (micropulse control). The CO305 carrier pcb is finishing-up and will be done early next week to be sent out for production (This will result in the full functionality of the ConOptics Model 305 synchronous count-down chassis to be reduced down to a single 6U euro card for micropulse selection.

Machine Protection System - The design for the new Machine Mode/Beam Mode VME boards is progressing very well. The I/O and addressing functions were finalized this week and the projected date for sending the board out for production is next week. Register assignments are

being passed on to the software group for EPICS screen construction. The documentation for the general purpose isolation module (GPIM64/05516-B-0009) has been prepared for our particular application in the new MPS system. The 64 bit VME isolation daughter board has been sent to EECAD for final checkplots.

The Fermi Lab trim card design was taken to a design engineering company for a quote on modifying the control card to hold a DSP. The quote will be ready next week. This has the potential to satisfy all of the FEL's DC Power needs from the 1 Amp trims to the 100 Amp inject and extract dipole strings. Work has begun on the 8 channels of floating power supplies for the arc quad windings. These will boost/buck fields in a larger magnet to produce the required quad fields. Installation of the DB/DJ correctors is in progress for the 2F and 3F region. Completed

cable removal for the old magnet shunts. This is the last large system targeted for cleanup/ removal to make room for new installation.

The pace has picked up for the new Gun HV interface. A system print is being finalized for getting the new HV tank wired into the safety systems and ready for operation. Chassis design and the VME Acromag fanout module design for the Gun HV interlocks and control continues as

well as work on the HV interlock PCB prototype.

MISC - some much needed maintenance was done on our data/web server (i.e. laser.jlab.org) The insertable dump water interlock requirements were identified and are being implemented by the mechanical support group.

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

The second GW underwent magnet field tests for core flatness on the primary magnet test stand. As with the first magnet, data indicates core field is near the plus or minus 1 part in 10000 flatness over the good field region.

MEI had a better measurement of the pole tip gap on the remaining 6 cores of the second batch of Magnets (for the UV line) at Diamond Tool, the subcontract machining house. Though the maximum spread in the readings is at the .003 inch level, (a hair thickness) our spec is one fifth of that. MEI is orchestrating the measurements and decisions on how the parts are to be brought into specification through selective grinding.

Injector Dipoles (DU/DV)

The measurements of the first GV continued. Repeatability is still a problem. We are working on hysteresis and power supply qualities.

We resolved our contract issues with Wang NMR. Because it is cheaper, the remaining magnets will be shipped by air freight.

The second GV and first GU arrived by air freight and look good.

Arc 180 Degree Dipoles (GY)

GY Coil Potting fixture is near completion (see photo).

We worked out the contractual issues associated with winding a fifth coil to replace the potentially unreliable first coil.

The conductor for the fifth coil is in insulation preparation.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

The GX cores have undergone gap inspection and are of similar high quality to the GUs.

At Wang NMR, potting fixtures are almost ready for the GQ and GX coils.

All GX and GQ coils are wound.

Quadrupoles

Trim Quad (QT)

Milhous shipped four trim quads with the auxiliary cooling plates on the side coils. We were ready to test, the stand worked for two minutes and then crashed. We are working on restoration.

Sextupole (SF)

DULY Research continued making substantial progress at finishing the Sextupole drawings. A third round of mark-ups (the coils) were sent back to them with minor corrections.

Octupole (OT)

The Octupole design is about half complete at DULY Research while they finish the Sextupole.

Beam Line and Vacuum

For the Arc Chamber job, Master Machine is continuing their work. They have subcontracted beveling to a Richmond firm.

Using the GW magnet, the test chambers were tested for their effect on core field and field flatness with the welds shaved. It removed an anomalous spike in one side of the 1/4 inch thick chambers.

The X Chamber fabrication is nearing completion.

The Jlab Shop has put fabrication of the four chambers for the optical chicane region on hold in order to complete more high priority work.

Design of girders for the regions around the wiggler was suspended for higher priority work.

The extraction "Y" chamber work was started in the Jlab Shop.

The stands for sextupoles and trim quads are due this week.

Girder and vacuum assembly continued

Low Conductivity Water piping installation continued in the FEL enclosure

The alignment crew continued stand placement.

WBS 11 (Optics):

This week's mechanical work on the optical cavity determined that the use of sprayed-on molybdenum disulfide (a JLab-approved UHV lubricant) onto a bare stainless steel leadscrew resulted in a ~ 2-3x reduction in drag torque when a Si/bronze nut was used. We plan to try sputtering this onto a leadscrew, and are having two silver-plated leadscrews stripped of silver so they can be plated. We are having parts fabricated to do the sputtering in-house, but are investigating whether it can be done by a coating vendor. We also will try tungsten disulfide.

Detailing continues on the optical transport turning cassette and insertable mirror designs. Our outgas rate test stand was rebuilt to test some optical cavity components.

We have been investigating what parameters influence the power and timing jitter of our ultrafast laser. A few additional parts for the FEL-FROG diagnostic were ordered. Parts for the injector-THz diagnostic are being fabricated. We inspected the mirrors that will be used in the injector light box and they handily met all optical specifications.

Other activities:

We worked on presentations to be given at the FEL conference and Boulder Damage Conferences. We hosted a visit by Lt. Fulvia Fiorani, NPS, and gave her materials and advice on the construction of a prototype active mirror stabilization system she is building for her Master's thesis. We spent some more time with potential vendors of the laser we'll use with the FEL external cavity (the pulse

stacker).